An object is attracted towards the origin with a force given by F = 6x^3 . a) What is the force at (1,0)? b) what is the work done by the when the object moves from (1,0) to (2,0)?


1
Expert's answer
2021-04-13T06:37:01-0400

a)

"F=6\cdot 1^3=6~N,"

b)

"A=\int _1^2\sqrt{1+F'^2(x)}dx=\int_1^2\sqrt{1+324x^4}dx\approx 42~J."